The height of an object dropped from a 205 feet tall building is given by h = -16t squared + 205. When will the object be 36 feet above the ground
Answer by ankor@dixie-net.com(22740)

The height of an object dropped from a 205 feet tall building is given by h = -16t squared + 205. When will the object be 36 feet above the ground
0 solutions
:
Substitute 36 for h:
-16t^2 + 205 = 36
:
-16t^2 = 36 - 205; Subtract 205 from both sides
:
16t^2 = 169; multiplied equation by -1
:
sqrt(16t^2) = sqrt(169); find the square root of both sides
:
4t = 13
:
t = 13/4
:
t = 3.25 seconds it will be 36 ft above the ground:
:
:
Check by substituting 3.25 for x, in -16t^2 + 205
-16(3.25^2) + 205 =
-16(10.5625) + 205
-169 + 205 = 36